OpenStudy (anonymous):

The cofunction of sin 25° is sec 65°. True or false? True False

OpenStudy (anonymous):

False sin 25º = 7.376 milliradians sec 65º = 41.3 milliradians

OpenStudy (campbell_st):

the cofunction is sin is \[\cos(90 - \theta) = \sin(\theta)\] you need to replace \[\theta... with... 25\]

OpenStudy (anonymous):

Thanks SO much!!! Fan ME!